Area of Operations: Wilmington Canyon, Hudson Canyon, Mid-Atlantic Bight, New Jersey Shelf, Georges Bank, Great South Channel, United States, North America, North Atlantic
Dates: July 06, 1977 to July 13, 1977
Chief scientist: Brad Butman
Objectives: 1. Recover 2 and deploy 4 bottom instrument packages as part of a continuing study of currents and sediment transport on the mid-Atlantic shelf and on Georges Bank. 2. Recover 1 current meter mooring and deploy 2 current meter moorings as part of the continuing study of currents and sediment transport. 3. Make underway XBT observations of water temperature structure and obtain profiles of temperature and conductivity (CTD) along selected transects running across shelf and along shelf. Measurements were designed to determine the position of the shelf-slope water front, to aid in interpretation of current meter and bottom tripod data, to map the distribution of cold pool water found on the shelf during summer, and to study mixing through Great South Channel. 4. Replace and relight as necessary surface marker floats at tripod locations. Floats serve to mark instrument locations and deter fishing activity.
Type of Activity: Mooring work; hydrographic survey;hydrographic survey
Information to be derived: Time series data; Samples and chemical analysis;
Summary: Preliminary data analysis: The current speed sensor on mooring #122 failed; otherwise data quality for 122 and 126 appears good. Data quality from mooring #118 is excellent. This mooring was deployed in Dec. 1976 and was in the water when the ARGO MERCHANT oil spill occurred. Current data should provide important information on possible fate of spilled oil. Crude anaylsis of XBT temperature observations aboard ship suggest mixing in Great South Channel. The temperature section off NJ showed a distinct cold pool. Hydrographic stations: 217 (XBT = 121, CTD = 88, surface salinity = 8). Moorings recovered: 3 (#118 deployed FA 76047 OCEANUS 17 Dec. 3-10, 1976; #122 and 126 both deployed FA 77009 WHITEFOOT Mar. 8 - Apr. 17, 1977). Moorings deployed: 6 (#128, 129, 130, 131, 132, and 133 (with transmissometer) all recovered FA 77032 ADVANCE II Sep. 7-23, 1977).
Contact: Brad Butman (firstname.lastname@example.org)
InfoBank: uses the identifier O-29-77-NA. Visit this site for more information.
- Download 77023.kml to view this field activity in Google Earth